OpenStudy (hba):

Adding (x+yi) in 6+2i will yield no number.The value of x and y is

OpenStudy (hba):

Does this mean i equate them equal to each other

OpenStudy (hba):

And get x=6 and y=2

OpenStudy (hba):

I think they are just trying to confuse

hartnn (hartnn):

add and equate to 0 , i think..

hartnn (hartnn):

it will definitaly yield a number

OpenStudy (hba):

Can you show me how will i do this

hartnn (hartnn):

adding 2 numbers always gives a number

OpenStudy (hba):

but 6+x=

OpenStudy (hba):

You are saying to put them equal to 0 so 6+x=0 x=-6

OpenStudy (hba):

Right ?

hartnn (hartnn):

i think its 'will yield no complex number' ?

OpenStudy (hba):

(y+2)i=0

OpenStudy (hba):

yi=-2i

OpenStudy (hba):

y=-2

OpenStudy (hba):

So x=-6 and y =-2

OpenStudy (hba):

@hartnn How do you interpret this ?

hartnn (hartnn):

yes, i think so

hartnn (hartnn):

means ?

OpenStudy (hba):

Did i do it the way you told me ?

hartnn (hartnn):

the question is weird. but i think that answers are correct

OpenStudy (hba):

Okay :) Thanks :)

hartnn (hartnn):

welcome :)

OpenStudy (hba):

@lgbasallote What do you think ?

OpenStudy (hba):

@VV-EFF Oeee Wasay ? Sahi Hai Ye

OpenStudy (hba):

x=-6 and y=-2 answer aya hai ?